sangWoAh:

What is the missing numerator? 2/7 = n/21 3 5 6 8

CocoCammron99:

How do you think we can eliminate

AZ:



B1:

The missing number is 6.

AZ:

we have to cross multiply

sangWoAh:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @B1 The missing number is 6. \(\color{#0cbb34}{\text{End of Quote}}\) oh ok

CocoCammron99:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @B1 The missing number is 6. \(\color{#0cbb34}{\text{End of Quote}}\) Can't give direct answers

ZombieLover:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @B1 The missing number is 6. \(\color{#0cbb34}{\text{End of Quote}}\) No Direct answers, please

sangWoAh:

so 2 times 3 is 6

sangWoAh:

and 7 times 3 is 21...

AZ:

Yes :)

sangWoAh:

OHHHHH I get ittt

B1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieLover \(\color{#0cbb34}{\text{Originally Posted by}}\) @B1 The missing number is 6. \(\color{#0cbb34}{\text{End of Quote}}\) No Direct answers, please \(\color{#0cbb34}{\text{End of Quote}}\) I am sorry. The answer is 6 because 21/3=7, meaning that the numbers were multiplied by 3. So you would multiply 2 by 3, making 6. And that is how I recieved my answer.

AZ:

But sometimes the numbers won't be this easy so if you follow what I posted above, you can always solve it!

ZombieLover:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @B1 \(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieLover \(\color{#0cbb34}{\text{Originally Posted by}}\) @B1 The missing number is 6. \(\color{#0cbb34}{\text{End of Quote}}\) No Direct answers, please \(\color{#0cbb34}{\text{End of Quote}}\) I am sorry. The answer is 6 because 21/3=7, meaning that the numbers were multiplied by 3. So you would multiply 2 by 3, making 6. And that is how I recieved my answer. \(\color{#0cbb34}{\text{End of Quote}}\) no worry make sure to explain next time :)

sangWoAh:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ But sometimes the numbers won't be this easy so if you follow what I posted above, you can always solve it! \(\color{#0cbb34}{\text{End of Quote}}\) ok!

AZ:

Here's an example 5/6 = x/720 We cross multiply 5 * 720 = 6 * x and we need to get x all by itself so we divide 6 on both sides and we get x = 5*720/6 x = 600

AZ:

Hopefully that helps you understand even more :)